OpenStudy (anonymous):

find the area of a regular pentagon with side equal to 3 and apothem equal to K A= ______K

Directrix (directrix):

The area of a regular pentagon is given by the formula (1/2) a p where a is the length of the apothem and p is the perimeter of the pentagon. A pentagon has 5 sides. Because it is regular, all sides have length 3. Hence, perimeter = 15. A = (1/2) k (15) A = (15/2) k or A = 7.5 k